OpenStudy (anonymous):

CAN SOMEONE PLEASE HELP ME? write an equation in point slope form of the line through point J (4,1) with slope -4

OpenStudy (johnweldon1993):

so point slope form is in the form \[\large y - y_o = m(x - x_0)\] where \(\large y_o = 1 \), \(\large m = -4\), \(\large x_o = 4\) so we have \[\large y - 1 = -4(x - 4)\] distribute the -4 \[\large y - 1 = -4x + 16\] Add 1 to both sides \[\large y = -4x + 17\] ^ would be your equation ^

undeadknight26 (undeadknight26):

:O i forgot all about this...thanks @johnweldon1993

OpenStudy (johnweldon1993):

Yeah no problem! Make sense @Christinabbbby ?

OpenStudy (anonymous):

yes, have time for a few more?

OpenStudy (johnweldon1993):

Sure :)

OpenStudy (anonymous):

write the equation for the horizontal line that contains point G (-8,8)

OpenStudy (johnweldon1993):

Alright so horizontal lines are in the form y = something... But whatever that y value is...it only passes through that same y-value... so if you have y = 1 ...the only y it will EVER pass through....will be 1 so if your point is (-8 , 8) the y-value here is 8... so y = 8

OpenStudy (anonymous):

Thank you so much, another one?

OpenStudy (johnweldon1993):

Go right ahead :)

OpenStudy (anonymous):

Which lines are parallel if m <4 =m <5 Justify your answer.

OpenStudy (anonymous):

Complete the two colum proof. GIVEN: <2 and <5 are supplementary PROVE: L/M

OpenStudy (anonymous):

there is another part but it wont let me post it, hold on